What is meant by the term temperature gradient?
1
Expert's answer
2020-05-28T13:42:28-0400

The term temperature gradient most commonly means how temperature changes with distance. For instance, we can say that the temperature gradient is -2.5°C per meter, which means that the temperature decreases 2.5°C with every meter away from some reference point, for instance, electric heater.

Mathematically, therefore, temperature gradient is a vector:

"\nabla T=\bigg(\frac{\partial T}{\partial x},\frac{\partial T}{\partial y},\frac{\partial T}{\partial z}\bigg)"